OpenStudy (anonymous):

Minimize 3x^2+4y^2, where x+y=7

OpenStudy (raden):

x+y=7 ---> y=7-x let k = 3x^2+4y^2 = 3x^2+4(7-x)^2 = 3x^2 + 4x^2-56x +196 = 7x^2-56x+196 k will be minimum when x=-b/2a